OpenStudy (anonymous):

what is the asymptote of this y=(x+1)/√(x^(2)+3)

OpenStudy (anonymous):

When do functions have asymptopes?

OpenStudy (campbell_st):

there are 2 horizontal asymptotes... they occur at \[y = \frac{x}{\pm x} \]

OpenStudy (anonymous):

The question is likely asking about vertical asymptotes. These are related to the restrictions to the x variable. In other words, you are looking for values of x that are not allowed.

OpenStudy (campbell_st):

well there are 2 types of asymptotes... vertical and oblique... your function, as written has no vertical asymptotes the reason is, no matter what value of x you have... squaring it will always give a postive number so adding a positive will still result in a positive and you can take the square root of any positive number... obliques... can be horizontal and can be a straight line in the form y = mx + b so if you are only being asked about vertical asymptotes, then you can say there are non.. and you now know why there are non...

OpenStudy (campbell_st):

or in math speak... the domain is all real x.

OpenStudy (campbell_st):

but there are restrictions in the range...

OpenStudy (campbell_st):

@mrleiss any comment...?

OpenStudy (anonymous):

You're right. I think most of the time the focus is on finding the VA. I hadn't looked at the HA at all.

OpenStudy (anonymous):

that is very helpful. i need the solution steps if possible so that i can understand more:) the thing is when given such question which i need to realtive max and min, i need to reate a variation table. but when it comes to sketching the graph using that variation table, i stuck. this is really dissapointing:( anyone.. help me, please?

OpenStudy (mertsj):

Since the degree of the numerator is smaller than the degree of the denominator, y = 0 is a horizontal asymptote.

OpenStudy (anonymous):

two vertical asymptotes at -3 and 3 and a horizontal asymptote

OpenStudy (anonymous):

there are 2 horizontal asymptotes. no vertical asymptote.

OpenStudy (campbell_st):

well done... thats i'm glad you are able to understand what I had written.

OpenStudy (agent0smith):

@mertsj that isn't really true, with the square root sign. To find the horizontal asymptotes, think about what happens as x approaches infinity.

OpenStudy (agent0smith):

\[\LARGE y=\frac{x+1} {\sqrt{x^2+3}}\] What happens to the numerator and denominator as x approaches infinity?

OpenStudy (anonymous):

but @campbell_st, i tried to create the variation table and compare with solution given, i cannot relate with the graph shown. if i were to refer the solution given, it is true we get 2 horizontal asymptotes but i still cannot get how...can you teach me?

OpenStudy (campbell_st):

well I looked at the problem this way as x gets extremely large the the fraction becomes \[\frac{x}{\sqrt{x^2}} = \frac{x}{-x}... and ....\frac{x}{x}\] which gives the horizontal asymptotes as y = -1 and y = 1

OpenStudy (anonymous):

can we assume like that?

OpenStudy (campbell_st):

well it works for me... thats how I found the asymptotes originally... not sure if they are correct or not. the problem is there are lots of ways to find asymptotes... I tend to use quick and dirty solutions... which get results... but thing about the method... as x gets extremely large... + 3 has no significance. and + 1 marginally less.

OpenStudy (agent0smith):

Yes, because as x becomes large, the 1 in x+1 becomes insignificant compared to the much larger x. And the 3 in x^2+3 becomes insignificant compared to the much larger x^2.

OpenStudy (anonymous):

ok noted. but have you guys @agent0smith, @campbell_st tried to sketch the graph using variation table? i'm not sure what you all call it but during my time it is called varition table..

OpenStudy (agent0smith):

Idk what that is, but i'd guess it's just a table of values. Try putting in very large positive values of x, and very small (negative) x values. or see this: https://www.google.com/search?q=y%3D(x%2B1)%2Fsqrt(x%5E(2)%2B3)&aq=f&oq=y%3D(x%2B1)%2Fsqrt(x%5E(2)%2B3)&aqs=chrome.0.57j0l3j62l2.5302j0&sourceid=chrome&ie=UTF-8 Zoom out and you'll see that y approaches 1 and -1 as x approaches +inf and -inf respectively.

OpenStudy (campbell_st):

well why not a table of values